nent CALCULATOoR ULL SCREEN Exercise 14-15 To pay for these bonds, Blossom had issued other bonds a month eartier bearing a lower interest rate. The newly issued bonds had a life of 10 years. The new bonds were issued at 101 (face value $312,000) Ignoring interest, compute the gain or loss. Loss on redermption Ignoring i Account Titles and Explanation Interest, record this refunding transaction. (If no entry is required, select No Entry for the account tities and enter o for the amounts. Credit account titles are y indented when amount is entered. Do not indent manually) Debit Credit (To record redemption of bonds payable) Version 4.24.11.5 All Rights Reserved. A Division of 2ohn iley &Sons,
Answer #1

Answer :-

outflow for bonds = 312000/100 *106 330720
unamortized discount 10000
Total 340720
inflow for the company = 312000/100*101 315120
Therefore loss on redemption = 25600
(1) Redemption of Bonds Payable )
Bonds outstanding a/c      Dr. 312000
Loss on Redemption          Dr. 18720
             To Bank a/c 330720
(2) Issuance of new bonds
Bank a/c                                 Dr. 315120
          To Bonds outstanding 312000
          To premium on bonds issued 3120